OpenStudy (anonymous):

Create your own division of polynomials problem. Demonstrate how this problem would be solved using both long division and synthetic division. I do not get this at all....

OpenStudy (anonymous):

well I did the same question as you and this is my work if you want to copy it I don't mind. 1) (x+2)^3 - 4 ---------------- = (x + 2)^2 with a remainder of -4. (x+2) ---------------------------------- 2) (x + 2) * (x + 3) * (x + 4) = x^3 + 9x^2 + 26x + 24 So let's divide x^3 + 9x^2 + 26x + 24 by (x + 2) Long division first: . . . . . .x^2 + 7x + 12 . . . . .---------------------------------------… x + 2 | x^3 + 9x^2 + 26x + 24 . . . . . x^3 + 2x^2 . . . . ----------------- . . . . . . . . . . 7x^2 + 26x . . . . . . . . . . .7x^2 + 14x + 24 . . . . . . . . . --------------------------- . . . . . . . . . . . . . . . . 12x + 24 . . . .. . . . . . . . . . . . . 12x + 24 . . . . . . . . . . . . . . . . ----------------- . . . . . . . . . . . . . . . . . . . . . . 0 <----------- remainder Since x^2 + 7x + 12 = (x + 3) * (x + 4) we know that this is correct. -------------------- Synthetic Division -2 | 1 9 26. 24. . . . . . . . . (because (x+2) = 0 implies x = -2 | . . -2 -14. -24 ------------------------------ . . .1. 7. 12. 0 This (1 7 12 0) is read as x^2 + 7x + 12 with a remainder of 0 Same answer as long division. Notice that I multiplied the 1 by (-2), the 7 by (-2) and the 12 by (-2)

OpenStudy (anonymous):

@LillieMax does this help

OpenStudy (anonymous):

Yes yes this does. Thanks so so much xD .

OpenStudy (anonymous):

yw